Together, research vessels A and B collected 500 seawater samples. If vessel B collected 56 more samples than vessel A, how many water samples did each vessel collect? Martha had 20 dimes and quarters worth $3.95. How many of each type of coin did she have? Find the 4th term of a geometric sequence whose first term is 6 and whose common ratio is 1/3. Find the 21st term of an arithmetic sequence whose first term is 9x and whose common difference is x + y.
Found 3 solutions by greenestamps, math_tutor2020, ikleyn:
Answer by greenestamps(13200)



These are problems on completely different subjects.

As the rules for this forum state, they should be submitted separately.


Answer by math_tutor2020(3817)


You should post one question at a time.
I'll focus on problem 1.

B = A+56
A+B = 500
A+A+56 = 500
2A+56 = 500
2A = 500-56
2A = 444
A = 444/2
A = 222
B = A+56
B = 222+56
B = 278

Answer: Vessel A collected 222 samples while vessel B collected 278 samples.

--------------------------------------------------------------------------

Problem 2 can be found here
https://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.1206400.html

I'll leave the others for the student to do.

Answer by ikleyn(52809)

.
(a) Together, research vessels A and B collected 500 seawater samples. If vessel B collected 56 more samples than vessel A, how many water samples did each vessel collect?
(b) Martha had 20 dimes and quarters worth $3.95. How many of each type of coin did she have?
(c) Find the 4th term of a geometric sequence whose first term is 6 and whose common ratio is 1/3.
(d) Find the 21st term of an arithmetic sequence whose first term is 9x and whose common difference is x + y.
~~~~~~~~~~~~~~~~~~~~~~~~


        All these problems were solved at this forum yesterday.
        Some of them were solved several times by different tutors.
        See the solutions under these links

(a)
https://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.1206402.html

(b)
https://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.1206400.html

(c)
https://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.1206401.html

(d)
https://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.1206399.html


///////////////////////


Please learn,  memorize and follow this simple rule :

        ONE  and  ONLY  ONE  problem per post.